# A New Impossibility Result for Online Bipartite Matching Problems

### v1

@@ 1 Introduction
Whether our upper bound is tight for some or all of the four problem variants mentioned above remains an open question. Notably, the bound involves the Euler’s number, in a manner analogous to the optimal competitive ratio of $1-\frac{1}{e}$ of the fully-adversarial Online Bipartite Matching problem. While we cannot determine whether our upper bound can be improved, we show that it possesses a certain extremal property and that it is the strongest bound possible for a natural class of user’s degrees distributions.
@@ 2 Overview
Our irregular cuckoo hashing instance contains $n$ users and $n$ ads, and assigns probability $\frac{1}{d\cdot(d-1)}$ to each user-degree $d\geq 2$ — in other words, it posits that the probability that a user has degree strictly larger than the generic positive integer $d$ is exactly $\nicefrac{{1}}{{d}}$. We obtained this distribution by greedily maximizing the number of users of degree $d=0,1,2,3,\ldots$ while guaranteeing that the random graph admits a quasi-complete matching from the user side — a matching that pairs all users except for at most $o(n)$ of them. Observe that, in a cuckoo hashing instance, the smaller the user-degrees the less likely it is for an online algorithm to match users. The goal of our distribution is to minimize the number of online matches while guaranteeing that the graph admits a quasi-complete matching — in other words, the distribution aims to minimize the numerator of the competitive ratio while making its denominator as large as possible. (We remark that most of the previous upper bounds were also based on graphs admitting quasi-complete matchings; see, e.g., [bk10, mos11, gm08, kvv90].)
@@ 5 Adaptation of the Analysis of Balister and Gerke to Unbounded Degree Distributions
In our arguments, we will make use of the following well-known concentration bounds (whose proofs can be found, e.g., in [dp09]): {fact}[Chernoff-Hoeffding inequality] Let $X_{1},\dots,X_{n}$ be independent random variables such that $X_{i}\in[a,b]$ almost surely. Let $X=\sum_{i=1}^{n}X_{i}$, then, for all $t>0$,
@@
$\Pr[|f(X_{1},\dots,X_{n})-\operatorname*{E}[f(X_{1},\dots,X_{n})]|>t]\leq 2\exp\left(-\frac{2\cdot t^{2}}{n\cdot d^{2}}\right).$
@@ 7 Extremality
Note that the function $g(y)=(1-y)\cdot\Phi(1-y,1,\Delta)$ is continuous and strictly decreasing for $y\in(0,1)$, moreover, $\lim_{y\rightarrow 0^{+}}g(y)=\infty$ and $g(1)=0$. Therefore, there exists a unique value of $y\in(0,1)$ that satisfies the equation and, by definition, such value is $\hat{w}_{1}$. While an exact computation of $\hat{w}_{1}$ would be challenging, we are only interested in small enough values of $\epsilon$ with respect to $\Delta$. We then compute $\hat{w}_{1}$ within an error term. Let $y_{1}=1-\Delta^{2}\epsilon+\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}$ and $y_{2}=y_{1}-\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}$. We have $y_{1},y_{2}\in(0,1)$ for a sufficiently small $\epsilon$, and,
$g(y_{1})$ $=\frac{\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}}{\Delta}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\right)^{2}}{\Delta+1}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\right)^{3}}{\Delta+2}\pm O(\epsilon^{4})$
$=\Delta\epsilon-\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}-\frac{2\Delta^{7}}{(\Delta+1)^{2}}\epsilon^{3}+\frac{\Delta^{6}}{\Delta+2}\epsilon^{3}\pm O(\epsilon^{4})$
$=\Delta\epsilon-\frac{\Delta^{6}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\cdot\epsilon^{3}\pm O(\epsilon^{4})<\Delta\epsilon,$
$g(y_{2})$ $=\frac{\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}}{\Delta}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}\right)^{2}}{\Delta+1}$
$\quad+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}\right)^{3}}{\Delta+2}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}\right)^{4}}{\Delta+3}\pm O(\epsilon^{5})$
$=\Delta\epsilon-\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{6}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}+\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{10}}{(\Delta+1)^{3}}\epsilon^{4}-\frac{2\Delta^{7}}{(\Delta+1)^{2}}\epsilon^{3}$
$\quad+\frac{2\Delta^{9}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{3}(\Delta+2)}\epsilon^{4}+\frac{\Delta^{6}}{\Delta+2}\epsilon^{3}-\frac{3\Delta^{9}}{(\Delta+1)(\Delta+2)}\epsilon^{4}+\frac{\Delta^{8}}{\Delta+3}\epsilon^{4}\pm O(\epsilon^{5})$
$=\Delta\epsilon+\frac{\Delta^{8}\left(\Delta^{4}+5\Delta^{3}+4\Delta^{2}-8\Delta+2\right)}{(\Delta+1)^{3}(\Delta+2)(\Delta+3)}\cdot\epsilon^{4}\pm O(\epsilon^{5})>\Delta\epsilon,$
@@
where the inequalities hold for a sufficiently small $\epsilon$ with respect to $\Delta$. Since $g(y)$ is continuous and decreasing, it must be $y_{2}<\hat{w}_{1}<y_{1}$, and therefore,
$\hat{w}_{1}=1-\Delta^{2}\epsilon+\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\pm O(\epsilon^{3}).$
@@
Moreover, by inverting the relation $\hat{w}_{1}=e^{-(1-w_{2})\mu}$ we get, $w_{2}=1+\frac{\ln\hat{w}_{1}}{\mu}$. Observe that $w_{2}$ is well-defined because, for a small enough $\epsilon$ with respect to $\Delta$, $y_{2}>e^{-\mu}$, and thus $\frac{\ln(\hat{w}_{1})}{\mu}>-1$.
@@
Note that $(1-\hat{w}_{1})^{\Delta-1}>0$ and moreover, we have, by using $\hat{w}_{1}=1-\Delta^{2}\epsilon+\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\pm O(\epsilon^{3})$,
$\epsilon(1+(\Delta-1)\hat{w}_{1})-\frac{1-\hat{w_{1}}}{\Delta}=\frac{\Delta^{2}}{\Delta+1}\cdot\epsilon^{2}\pm O(\epsilon^{3}).$
@@
Thus, for a small enough $\epsilon$ with respect to $\Delta$, we have $\alpha<1-\frac{1}{\Delta}+\epsilon$, in particular,
$\alpha$ $=1-\frac{1}{\Delta}+\epsilon-\left(\Delta^{2}\cdot\epsilon-\frac{\Delta^{5}}{\Delta+1}\cdot\epsilon^{2}\pm O(\epsilon^{3})\right)^{\Delta-1}\cdot\left(\frac{\Delta^{2}}{\Delta+1}\cdot\epsilon^{2}\pm O(\epsilon^{3})\right)$
$=1-\frac{1}{\Delta}+\epsilon-\left(\Delta^{2\Delta-2}\cdot\epsilon^{\Delta-1}\pm O(\epsilon^{\Delta})\right)\cdot\left(\frac{\Delta^{2}}{\Delta+1}\cdot\epsilon^{2}\pm O(\epsilon^{3})\right)$
$=1-\frac{1}{\Delta}+\epsilon-\frac{\Delta^{2\Delta}}{\Delta+1}\cdot\epsilon^{\Delta+1}\pm O(\epsilon^{\Delta+2}).$
@@
We can now use the previous Lemma to prove the main result of this Section. {theoremrep} Suppose that there exists a constant $i\geq 1$ such that $\Pr[P\leq j]=1-\frac{1}{j}$ for each $j\in\{1,2,\ldots,i-1\}$, and $\Pr[P\leq i]>1-\frac{1}{i}$. Then, with probability $1-o(1)$, the instance $\operatorname*{ICH}(n,n,P)$ admits no matching that matches at least $\Pr[P\leq i]\cdot n-o(n)$ users of degree $\leq i$. {appendixproof} Consider first the case $i=1$. Note that, if $\Pr[P=0]>0$, then, from a simple concentration bound, we have that, with probability $1-o(1)$, at least $\Pr[P=0]\cdot n-o(n)$ users of degree $\leq i$ cannot be matched. Consider now the case $\Pr[P=0]=0$ and $\Pr[P=1]=\epsilon>0$. With probability $1-o(1)$, there are at least $(\epsilon-o(1))n$ users of degree 1. Let $S$ be the set of these users. Consider the generic ad $a$. We compute the probabilities that this ad has 00 or $1$ neighbors in $S$: $\Pr[a\text{ has no neighbor in }S]=\left(1-\frac{1}{n}\right)^{|S|}=e^{-\epsilon}\pm o(1)$, and $\Pr[a\text{ has exactly 1 neighbor in }S]=|S|\cdot\frac{1}{n}\cdot\left(1-\frac{1}{n}\right)^{|S|-1}=\epsilon\cdot e^{-\epsilon}\pm o(1)$. Thus, if we let $T$ be the set of ads that have $2$ or more neighbors in $S$, we have that, with probability $1-o(1)$, $|T|\geq(1-(\epsilon+1)\cdot e^{-\epsilon})\cdot n-o(n)$. Given that each user in $S$ has degree $1$, in each matching $M$, each ad in $T$ contributes to the count of users unmatched by $M$ with (at least) a unique user. Therefore, at least $(1-(\epsilon+1)\cdot e^{-\epsilon})\cdot n-o(n)$ users of degree $1$ will remain unmatched.

### v2

@@ 1 Introduction
Whether our upper bound is tight for some or all of the four problem variants mentioned above remains an open question. Notably, the bound involves the Euler’s number, in a manner analogous to the optimal competitive ratio of $1-\frac{1}{e}$ of the fully-adversarial Online Bipartite Matching problem. While we cannot determine whether our upper bound can be improved, we show that it possesses a certain extremal property and that it is the strongest bound possible for a natural class of instances.
@@ 2 Overview
Our irregular cuckoo hashing instance contains $n$ users and $n$ ads, and assigns probability $\frac{1}{d\cdot(d-1)}$ to each user-degree $d\geq 2$ — in other words, it posits that the probability that a user has degree strictly larger than the generic positive integer $d$ is exactly $\nicefrac{{1}}{{d}}$. We obtained this distribution by greedily maximizing the number of users of degree $d=0,1,2,3,\ldots$ while guaranteeing that the random graph admits a quasi-complete matching from the user side — a matching that pairs all users except for at most $o(n)$ of them. Observe that, in a cuckoo hashing instance, the smaller the user-degrees the less likely it is for an online algorithm to match users. The goal of our distribution is to minimize the number of online matches while guaranteeing that the graph admits a quasi-complete matching — in other words, the distribution aims to minimize the numerator of the competitive ratio while making its denominator as large as possible. (We remark that most previous upper bounds were also based on graphs admitting quasi-complete matchings; see, e.g., [bk10, mos11, gm08, kvv90].)
@@ 5 Adaptation of the Analysis of Balister and Gerke to Unbounded Degree Distributions
In our arguments, we will make use of the following well-known concentration bounds (whose proofs can be found, e.g., in [dp09]): {fact}[Chernoff-Hoeffding’s inequality] Let $X_{1},\dots,X_{n}$ be independent random variables such that $X_{i}\in[a,b]$ almost surely. Let $X=\sum_{i=1}^{n}X_{i}$, then, for all $t>0$,
@@
$\Pr[|f(X_{1},\dots,X_{n})-\operatorname*{E}[f(X_{1},\dots,X_{n})]|\geq t]\leq 2\exp\left(-\frac{2\cdot t^{2}}{n\cdot d^{2}}\right).$
@@ 7 Extremality
Note that the function $g(y)=(1-y)\cdot\Phi(1-y,1,\Delta)$ is continuous and strictly decreasing for $y\in(0,1)$, moreover, $\lim_{y\rightarrow 0^{+}}g(y)=\infty$ and $g(1)=0$. Therefore, there exists a unique value of $y\in(0,1)$ that satisfies the equation and, by definition, such value is $\hat{w}_{1}$. While an exact computation of $\hat{w}_{1}$ would be challenging, we are only interested in small enough values of $\epsilon$ with respect to $\Delta$. We then compute $\hat{w}_{1}$ within an error term. Let $y_{1}=1-\Delta^{2}\epsilon+\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}$ and $y_{2}=y_{1}-\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}$. We have $y_{1},y_{2}\in(0,1)$ for a sufficiently small $\epsilon$. In particular, we will select $\epsilon<\frac{1}{2\cdot 10^{3}\cdot(3\Delta)^{7\Delta}}$. Note that $y_{1}\geq y_{2}\geq\nicefrac{{1}}{{2}}$. Moreover,
$\sum_{j=3}^{\infty}\frac{(1-y_{1})^{j+1}}{j+\Delta}$ $\leq\sum_{j=3}^{\infty}(1-y_{1})^{j+1}=(1-y_{1})\left(\sum_{j=0}^{\infty}(1-y_{1})^{j}-\sum_{j=0}^{2}(1-y_{1})^{j}\right)$
$=(1-y_{1})\left(\frac{1}{y_{1}}-\frac{1-(1-y_{1})^{3}}{y_{1}}\right)=\frac{(1-y_{1})^{4}}{y_{1}}$
$\leq 2(1-y_{1})^{4}=2\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\right)^{4}\leq 2\left(\Delta^{2}\epsilon\right)^{4}\leq 2\Delta^{8}\epsilon^{4},$
since $\Delta^{2}\epsilon>\Delta^{4}\epsilon^{2}\geq\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}$ for the chosen $\epsilon$. By using that $\frac{\Delta^{6}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\geq\frac{\Delta^{5}}{12}$ for $\Delta\geq 1$, we have,
$g(y_{1})$ $=\frac{\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}}{\Delta}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\right)^{2}}{\Delta+1}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\right)^{3}}{\Delta+2}+\sum_{j=3}^{\infty}\frac{(1-y_{1})^{j+1}}{j+\Delta}$
$\leq\Delta\epsilon-\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\right)^{2}}{\Delta+1}+\frac{\Delta^{6}\epsilon^{3}}{\Delta+2}+2\Delta^{8}\epsilon^{4}$
$\leq\Delta\epsilon-\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}-\frac{2\Delta^{7}}{(\Delta+1)^{2}}\epsilon^{3}+\frac{\Delta^{10}}{(\Delta+1)^{3}}\epsilon^{4}+\frac{\Delta^{6}}{\Delta+2}\epsilon^{3}+2\Delta^{8}\epsilon^{4}$
$\leq\Delta\epsilon-\frac{\Delta^{6}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\cdot\epsilon^{3}+3\Delta^{8}\epsilon^{4}$
$\leq\Delta\epsilon-\frac{\Delta^{5}}{12}\cdot\epsilon^{3}+3\Delta^{8}\epsilon^{4}$
$<\Delta\epsilon,$
where the last inequality holds because $\epsilon<\frac{1}{36\cdot\Delta^{3}}$. Now, observe that $\frac{2\Delta^{12}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{3}(\Delta+2)}\leq 6\Delta^{10}$ and $\frac{\Delta^{8}\left(\Delta^{4}+5\Delta^{3}+4\Delta^{2}-8\Delta+2\right)}{(\Delta+1)^{3}(\Delta+2)(\Delta+3)}\geq\frac{\Delta^{7}}{96}$ for $\Delta\geq 1$. Note also that $(1-x)^{4}\geq(1-8x)$ for $x\in(0,1/8)$. We lower bound $g(y_{2})$ by truncating the sum to the fourth term,
$g(y_{2})$ $>\frac{\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}}{\Delta}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}\right)^{2}}{\Delta+1}$
$\quad+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}\right)^{3}}{\Delta+2}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}\right)^{4}}{\Delta+3}$
$\geq\Delta\epsilon-\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{6}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{7}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}\right)^{2}}{\Delta+1}$
$\quad+\frac{\left(\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}\right)^{3}}{\Delta+2}+\frac{\left(\Delta^{2}\epsilon-\Delta^{4}\epsilon^{2}\right)^{4}}{\Delta+3}$
$\geq\Delta\epsilon-\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{6}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{2}(\Delta+2)}\epsilon^{3}+\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}+\frac{\Delta^{10}}{(\Delta+1)^{3}}\epsilon^{4}-\frac{2\Delta^{7}}{(\Delta+1)^{2}}\epsilon^{3}$
$\quad+\frac{2\Delta^{9}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{3}(\Delta+2)}\epsilon^{4}-\frac{2\Delta^{12}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{3}(\Delta+2)}\epsilon^{5}$
$\quad+\frac{\Delta^{6}}{\Delta+2}\epsilon^{3}-\frac{3\Delta^{9}}{(\Delta+1)(\Delta+2)}\epsilon^{4}-\frac{\Delta^{15}}{(\Delta+1)^{3}(\Delta+2)}\epsilon^{6}+\frac{(\Delta^{2}\epsilon)^{4}(1-\Delta^{2}\epsilon)^{4}}{\Delta+3}$
$\geq\Delta\epsilon+\frac{\Delta^{10}}{(\Delta+1)^{3}}\epsilon^{4}+\frac{2\Delta^{9}(\Delta^{2}+2\Delta-1)}{(\Delta+1)^{3}(\Delta+2)}\epsilon^{4}-\frac{3\Delta^{9}}{(\Delta+1)(\Delta+2)}\epsilon^{4}+\frac{\Delta^{8}\epsilon^{4}(1-8\Delta^{2}\epsilon)}{\Delta+3}-7\Delta^{10}\epsilon^{5}$
$\geq\Delta\epsilon+\frac{\Delta^{8}\left(\Delta^{4}+5\Delta^{3}+4\Delta^{2}-8\Delta+2\right)}{(\Delta+1)^{3}(\Delta+2)(\Delta+3)}\cdot\epsilon^{4}-15\Delta^{10}\epsilon^{5}$
$\geq\Delta\epsilon+\frac{\Delta^{7}}{96}\epsilon^{4}-15\Delta^{10}\epsilon^{5}$
$>\Delta\epsilon,$
@@
where the last inequality follows since $\epsilon<\frac{1}{2\cdot 10^{3}\cdot\Delta^{3}}$. Since $g(y)$ is continuous and decreasing, it must be $y_{2}<\hat{w}_{1}<y_{1}$, and therefore,
$\hat{w}_{1}=1-\Delta^{2}\epsilon+\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}-R(\Delta,\epsilon),$
@@
where $R(\Delta,\epsilon)=y_{1}-\hat{w}_{1}\leq y_{1}-y_{2}\leq 3\Delta^{6}\epsilon^{3}$. Moreover, by inverting the relation $\hat{w}_{1}=e^{-(1-w_{2})\mu}$ we get, $w_{2}=1+\frac{\ln\hat{w}_{1}}{\mu}$. Observe that $w_{2}$ is well-defined because $y_{2}\geq 1/2>e^{-\mu}$, and thus $\frac{\ln(\hat{w}_{1})}{\mu}>-1$.
@@
Note that $(1-\hat{w}_{1})^{\Delta-1}>0$ and moreover, we have, by using $\hat{w}_{1}=1-\Delta^{2}\epsilon+\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}-R(\Delta,\epsilon)$,
$\epsilon(1+(\Delta-1)\hat{w}_{1})-\frac{1-\hat{w_{1}}}{\Delta}$
$=\epsilon\left(1+(\Delta-1)\left(1-\Delta^{2}\epsilon+\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}-R(\Delta,\epsilon)\right)\right)-\frac{\Delta^{2}\epsilon-\frac{\Delta^{5}}{\Delta+1}\epsilon^{2}+R(\Delta,\epsilon)}{\Delta}$
$=\epsilon\left(\Delta-(\Delta-1)\Delta^{2}\epsilon+\frac{\Delta^{5}(\Delta-1)}{\Delta+1}\epsilon^{2}-(\Delta-1)R(\Delta,\epsilon)\right)-\Delta\epsilon+\frac{\Delta^{4}}{\Delta+1}\epsilon^{2}-\frac{R(\Delta,\epsilon)}{\Delta}$
$=\frac{\Delta^{2}}{\Delta+1}\cdot\epsilon^{2}-R_{2}(\Delta,\epsilon),$
@@
with $R_{2}(\Delta,\epsilon)=\frac{R(\Delta,\epsilon)}{\Delta}+\epsilon(\Delta-1)R(\Delta,\epsilon)-\frac{\Delta^{5}(\Delta-1)}{(\Delta+1)}\epsilon^{3}$ and therefore, $|R_{2}(\Delta,\epsilon)|\leq 7\Delta^{5}\epsilon^{3}$. Thus, since $\epsilon<\frac{1}{7\Delta^{3}(\Delta+1)}$, we have $\alpha<1-\frac{1}{\Delta}+\epsilon$. In particular,
$\alpha$ $=1-\frac{1}{\Delta}+\epsilon-\left(\Delta^{2}\cdot\epsilon-\frac{\Delta^{5}}{\Delta+1}\cdot\epsilon^{2}+R(\Delta,\epsilon)\right)^{\Delta-1}\cdot\left(\frac{\Delta^{2}}{\Delta+1}\cdot\epsilon^{2}-R_{2}(\Delta,\epsilon)\right)$
$=1-\frac{1}{\Delta}+\epsilon-\left(\Delta^{2\Delta-2}\cdot\epsilon^{\Delta-1}+R_{3}(\Delta,\epsilon)\right)\cdot\left(\frac{\Delta^{2}}{\Delta+1}\cdot\epsilon^{2}-R_{2}(\Delta,\epsilon)\right)$
$=1-\frac{1}{\Delta}+\epsilon-\frac{\Delta^{2\Delta}}{\Delta+1}\cdot\epsilon^{\Delta+1}+R_{4}(\Delta,\epsilon),$
where $R_{3}(\Delta,\epsilon)$ is the sum of $13^{\Delta-1}-$ 1 terms, each no larger than $(3\Delta)^{6(\Delta-1)}\epsilon^{\Delta}$ in absolute value, and thus, $|R_{3}(\Delta,\epsilon)|\leq(3\Delta)^{7(\Delta-1)}\epsilon^{\Delta}$. Moreover,
$|R_{4}(\Delta,\epsilon)|$ $\leq\left|\frac{\Delta^{2}\epsilon^{2}R_{3}(\Delta,\epsilon)}{\Delta+1}\right|+\left|R_{3}(\Delta,\epsilon)R_{2}(\Delta,\epsilon)\right|+\left|R_{2}(\Delta,\epsilon)\Delta^{2\Delta-2}\epsilon^{\Delta-1}\right|$
$\leq(3\Delta)^{7\Delta-5}\epsilon^{\Delta+2}+7(3\Delta)^{7\Delta-2}\epsilon^{\Delta+3}+7\Delta^{2\Delta+3}\epsilon^{\Delta+2}$
$\leq 15\cdot(3\Delta)^{7\Delta}\cdot\epsilon^{\Delta+2}.$
@@
Thus, $\psi_{\Delta}(\epsilon)=\frac{\Delta^{2\Delta}}{\Delta+1}\epsilon^{\Delta+1}-R_{4}(\Delta,\epsilon)>0$ since $\epsilon<\frac{1}{15\cdot(3\Delta)^{7\Delta}}$. We can now use the previous Lemma to prove the main result of this Section. {theoremrep} Suppose that there exists a constant $i\geq 1$ such that $\Pr[P\leq j]=1-\frac{1}{j}$ for each $j\in\{1,2,\ldots,i-1\}$, and $\Pr[P\leq i]>1-\frac{1}{i}$. Then, with probability $1-o(1)$, the instance $\operatorname*{ICH}(n,n,P)$ admits no matching that matches at least $\Pr[P\leq i]\cdot n-o(n)$ users of degree $\leq i$. {appendixproof} Consider first the case $i=1$. Note that, if $\Pr[P=0]>0$, then, from a simple concentration bound, we have that, with probability $1-o(1)$, at least $\Pr[P=0]\cdot n-o(n)$ users of degree $\leq i$ cannot be matched. Consider now the case $\Pr[P=0]=0$ and $\Pr[P=1]=\epsilon>0$. With probability $1-o(1)$, there are at least $(\epsilon-o(1))n$ users of degree 1. Let $S$ be the set of these users. Consider the generic ad $a$. We compute the probabilities that this ad has 00 or $1$ neighbors in $S$: $\Pr[a\text{ has no neighbor in }S]=\left(1-\frac{1}{n}\right)^{|S|}=e^{-\epsilon}\pm o(1)$, and $\Pr[a\text{ has exactly 1 neighbor in }S]=|S|\cdot\frac{1}{n}\cdot\left(1-\frac{1}{n}\right)^{|S|-1}=\epsilon\cdot e^{-\epsilon}\pm o(1)$. Thus, if we let $T$ be the set of ads that have $2$ or more neighbors in $S$, we have that, with probability $1-o(1)$, $|T|\geq(1-(\epsilon+1)\cdot e^{-\epsilon})\cdot n-o(n)$. Given that each user in $S$ has degree $1$, in each matching $M$, each ad in $T$ contributes to the count of users unmatched by $M$ with (at least) a unique user. Therefore, at least $(1-(\epsilon+1)\cdot e^{-\epsilon})\cdot n-o(n)$ users of degree $1$ will remain unmatched.